OpenStudy (anonymous):

Help me please can someone help me I've been on here all day needing help with this problem ? Your mortgage payment is $954.32/month. Using proportions, what is the minimum yearly amount you must have in realized income to keep your housing expenses in the acceptable range?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[\frac{ $954.32 }{ 1 month }=\frac{ X }{ 12 months }\]

OpenStudy (anonymous):

so i divide 954.32/12 then by 1

OpenStudy (anonymous):

\[954.32\times 12=X \times 1\]

OpenStudy (anonymous):

X= $954.32*12

OpenStudy (anonymous):

11451.84

OpenStudy (anonymous):

what about You have a mortgage payment of $1,248.90/month. Using proportions, what is the minimum amount you must have in realized income per month to keep your housing expense in the acceptable range?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

did you ever get the answer to the last one?

OpenStudy (anonymous):

@WeepingAngel123